# Fino1: On the Transferability of Reasoning-Enhanced LLMs to Finance

### v2

@@ A.2 LLMs in Financial Applications
The application of LLMs in finance has gained significant attention, with models being adapted for tasks such as financial text analysis Yang et al. (2020), market sentiment prediction Yang et al. (2023), and automated trading strategies Kou et al. (2024). Recent work has focused on developing finance-specific models like BloombergGPT Wu et al. (2023), FinGPT Yang et al. (2023) and PIXIU Xie et al. (2023), OpenFinLLMs Xie et al. (2024b), which are pretrained on financial corpora to better understand domain-specific terminology and concepts. These models have demonstrated effectiveness in tasks such as financial sentiment analysis, earnings call analysis, and regulatory compliance checking.
A.3 Reasoning Capabilities in Financial Tasks
Financial reasoning presents unique challenges that combine numerical computation, domain knowledge, and logical inference. Recent studies have explored LLMs’ capabilities in financial reasoning tasks, including numerical reasoning over financial documents Chen et al. (2021), mathematical problem-solving in financial contexts Srivastava et al. (2024), and structured data interpretation Liu et al. (2025a). These works highlight the importance of both domain expertise and reasoning abilities in financial applications. Our work extends these research directions by systematically evaluating how reasoning enhancements in state-of-the-art LLMs translate to financial domain tasks, providing insights into the effectiveness of different reasoning strategies and identifying areas for improvement in financial reasoning capabilities.
Appendix B Details Data Source
@@
  • FinQA Chen et al. (2022): FinQA consists of over 5,000 open-ended question-answer pairs derived from financial reports and tables. It integrates both textual and tabular data, covering numerical reasoning tasks such as arithmetic operations, comparison, and multi-step calculations, making it well-suited for training reasoning-enhanced models.
@@
  • DocFinQA Reddy et al. (2024b): An extension of FinQA designed for long-context reasoning. It pairs FinQA-style questions with full-length financial filings, increasing the average context length from under 700 words to approximately 123,000 words. This dataset better simulates real-world document scale and complexity, enabling robust evaluation of retrieval and reasoning over long financial documents.
@@
  • TAT-QA Zhu et al. (2021): A large-scale financial QA dataset created from real financial reports. It combines both textual and tabular data and emphasizes diverse numerical reasoning tasks, including arithmetic operations, counting, sorting, and comparison. TAT-QA reflects real-world challenges in financial document interpretation.
@@
  • DocMath-Eval Zhao et al. (2024b): This benchmark suite evaluates mathematical reasoning in domain-specific contexts. We utilize two subsets:
@@
    • DocMath-CompLong: Constructed from quarterly and annual reports, it involves reasoning across long documents and multiple tables, presenting a challenging multi-step numerical reasoning setting.
@@
    • DocMath-SimpShort: Built from FinQA and TAT-QA, this subset focuses on localized numerical reasoning tasks over short documents containing a single table.
@@
  • BizBench-QA Krumdick et al. (2024): Contains 14,400 financial QA pairs across multiple task types, including SEC-NUM, CodeFinQA, CodeTAT-QA, and FinCode. These tasks span code generation, numerical span identification, and financial mathematical reasoning from various perspectives, enriching the training corpus with broader financial task coverage.
@@
Appendix C Error Analysis
@@
Limitations
C.1 Future Direction
Ethics Statement
Appendix D Prompt

### v3

@@ A.2 LLMs in Financial Applications
The application of LLMs in finance has gained significant attention, with models being adapted for tasks such as financial text analysis [42], market sentiment prediction [41], and automated trading strategies [13]. Recent work has focused on developing finance-specific models like BloombergGPT [36], FinGPT [41] and PIXIU [38], OpenFinLLMs [39], Plutus[24], which are pretrained on financial corpora to better understand domain-specific terminology and concepts. These models have demonstrated effectiveness in tasks such as financial sentiment analysis, earnings call analysis, and regulatory compliance checking.
A.3 Financial Reasoning
Financial reasoning presents unique challenges that combine numerical computation, domain knowledge, and logical inference. Recent studies have explored LLMs’ capabilities in financial reasoning tasks, including numerical reasoning over financial documents [4], mathematical problem-solving in financial contexts [29], and structured data interpretation [18]. These works highlight the importance of both domain expertise and reasoning abilities in financial applications. Our work extends these research directions by systematically evaluating how reasoning enhancements in state-of-the-art LLMs translate to financial domain tasks, providing insights into the effectiveness of different reasoning strategies and identifying areas for improvement in financial reasoning capabilities. Additionally, we focus on benchmarking these models across diverse financial datasets, emphasizing multi-table reasoning, long-context comprehension, and quantitative analysis. This comprehensive evaluation not only sheds light on current model limitations but also informs future research directions for developing more robust financial reasoning models.
Appendix B Details of Data Source
B.1 Data source examples for other works
This section present some examples of the data from other models. In convince of reviewing, we have translated all the original context into English if they are not English.
Finance-Instruct Explain tradeoffs between fiscal and monetary policy as tools in a nations economic toolkit. Provide examples of past instances when each were utilized, the economic conditions that led to them being deployed, their intended effects, and an evaluation of their relative efficacy and consequences.
CFLUE Assume you are a financial industry expert. Please answer the following question. Note: The question is a single-choice question. Only return the most appropriate option. If multiple options are suitable, only return the most accurate one. Among the following options, which financial institution is not permitted to participate in the interbank lending market? A. Clearing Center B. Insurance Company C. Securities Company D. Fund Management Company
CFLUE Is the sentiment of this text positive, negative, or neutral? Feeling exhausted from averaging down.
Finance-IQ When considering the pricing of life insurance products, several key factors come into play, including market forces, business operating costs, and government regulations. Among these factors, which one primarily determines the minimum limit of life insurance product prices? A. Market forces and business operating costs B. Business operating costs C. Government regulations D. Market forces
B.2 Data source for FinCoT
@@
  • FinQA [5]: FinQA consists of over 5,000 open-ended question-answer pairs derived from financial reports and tables. It integrates both textual and tabular data, covering numerical reasoning tasks such as arithmetic operations, comparison, and multi-step calculations, making it well-suited for training reasoning-enhanced models.
@@
  • DocFinQA [27]: An extension of FinQA designed for long-context reasoning. It pairs FinQA-style questions with full-length financial filings, increasing the average context length from under 700 words to approximately 123,000 words. This dataset better simulates real-world document scale and complexity, enabling robust evaluation of retrieval and reasoning over long financial documents.
@@
  • TAT-QA [47]: A large-scale financial QA dataset created from real financial reports. It combines both textual and tabular data and emphasizes diverse numerical reasoning tasks, including arithmetic operations, counting, sorting, and comparison. TAT-QA reflects real-world challenges in financial document interpretation.
  • ConvFinQA [6]: A financial question answering dataset designed to handle multi-turn conversations involving financial tables and text. ConvFinQA builds upon FinQA by incorporating follow-up questions and context retention across multiple turns, simulating more realistic financial analysis scenarios. It provides expert-annotated sub-questions that decompose complex financial questions into granular steps, facilitating the generation of structured reasoning paths and promoting systematic financial reasoning.
@@
  • DocMath-Eval [45]: This benchmark suite evaluates mathematical reasoning in domain-specific contexts. We utilize two subsets:
@@
    • DocMath-CompShort[45]: Constructed from quarterly and annual reports, it involves reasoning across long documents and multiple tables, presenting a challenging multi-step numerical reasoning setting.
@@
    • DocMath-SimpShort[45]: Built from FinQA and TAT-QA, this subset focuses on localized numerical reasoning tasks over short documents containing a single table.
  • Econ_Logic [25]: It consists of financial logic reasoning tasks that help the model to further understand and have more insights of financial terms and knowledge.
@@
  • BizBench-QA [14]: Contains financial QA pairs across multiple task types, including SEC-NUM, CodeFinQA, CodeTAT-QA, and FinCode. These tasks span code generation, numerical span identification, and financial mathematical reasoning from various perspectives, enriching the training corpus with broader financial task coverage.
FinQA Please answer the given financial question based on the context. Context: measurement point december 31 the priceline group nasdaq composite index s&p 500 rdg internet composite . |measurement pointdecember 31|the priceline group inc .|nasdaqcomposite index|s&p 500index|rdg internetcomposite| |2010|100.00|100.00|100.00|100.00| |2011|117.06|100.53|102.11|102.11| |2012|155.27|116.92|118.45|122.23| |2013|290.93|166.19|156.82|199.42| |2014|285.37|188.78|178.29|195.42| |2015|319.10|199.95|180.75|267.25|. Question: what was the percent of the growth of the the priceline group inc . from 2014 to 2015 Answer:
Appendix C Example of CoT data curation
Here is an example of applying the guideline and iterative refinement process to generate a CoT reasoning path based on the given context and question.
Figure 3: Example of iterative refinement of the CoT process with a guidance.
BizBench From the results reflected in the following pro forma information because of normal production declines, changes in commodity prices, future acquisitions and divestitures, future development and exploration activities and other factors. | | | | | | | | | | | | | --- | --- | --- | --- | --- | --- | --- | --- | --- | --- | --- | | | | | | | | | | | | | | | | | | | | (In millions, except per share amounts) | | Six Months Ended June 30, 2021 | | Pro forma revenue | | \$ | 1,802 | | | | | Pro forma net income | | 71 | | | | | | | | | | | | | | | | | | | | | | | | | | | | | Pro forma basic earnings per share | | \$ | 0.09 | | | | | Pro forma diluted earnings per share | | \$ | 0.09 | | | | 3. Properties and Equipment, NetProperties and equipment, net are comprised of the following: | | | | | | | | | | | | | | | | | --- | --- | --- | --- | --- | --- | --- | --- | --- | --- | --- | --- | --- | --- | --- | | | | | | | | | | | | | | | | | | (In millions) | | June 30,2022 | | December 31,2021 | | Proved oil and gas properties | | \$ | 16,102 | | | \$ | 15,340 | | | Unproved oil and gas properties | | 5,292 | | | 5,316 | | | Gathering and pipeline systems | | 423 | | | 395 | | | Land, buildings and other equipment | | 144 | | | 140 | | | Finance lease right-of-use asset | | 24 | | | 20 | | | | | 21,985 | | | 21,211 | | | Accumulated depreciation, depletion and amortization | | (4,578) | | | (3,836) | | | | | \$ | 17,407 | | | \$ | 17,375 | | As of June 30, 2022, the Company had no projects with exploratory well costs capitalized for more than one year after drilling. | | | | | | | | | | | | | | | | | --- | --- | --- | --- | --- | --- | --- | --- | --- | --- | --- | --- | --- | --- | --- | | | | | | | | | | | | | | | | | | (In millions) | | June 30,2022 | | December 31,2021 | | 6.51% weighted-average private placement senior notes (1) | | \$ | 37 | | | \$ | 37 | | | 5.58% weighted-average private placement senior notes (2) | | 87 | | | 87 | | | 3.65% weighted-average private placement senior notes | | 825 | | | 825 | | | 4.375% senior notes due June 1, 2024 | | 750 | | | 750 | | | 3.90% senior notes due May 15, 2027 | | 750 | | | 750 | | | 4.375% senior notes due March 15, 2029 | | 500 | | | 500 | | | Revolving credit facility | | - | | | - | | | Net premium (discount) | | 164 | | | 185 | | | Unamortized debt issuance costs | | (8) | | | (9) | | | | | \$ | 3,105 | | | \$ | 3,125 | | (1) Includes \$37 million of current portion of long-term debt at June 30, 2022.(2) Includes \$87 million of current portion of long-term debt at June 30, 2022.At June 30, 2022, the Company was in compliance with all financial and other covenants for both its revolving credit facility and senior notes. Question: What is the pro forma net income?
Appendix D Details of experiments
D.1 Model training
We list the details of the training process here. The SFT process is configured with a learning rate of 5e-6 and a weight decay of 0.1. Training is conducted for 3 epoch with a batch size of 2 per GPU and a gradient accumulation of 2 steps, effectively resulting in an effective batch size of 2. The gradient checkpointing is enabled to optimize memory usage. For RL, the training process is configured with a learning rate specified set to 5e-6. The gradient accumulation is set to 2 steps, and a gradient accumulation of 2 steps, effectively resulting in an effective batch size based on the configured gradient_accumulation_steps. The training is conducted using 3 GPUs, each handling a separate training process. And a total of 500 steps is used to train the model.
D.2 Model evaluation
For large models (e.g., DeepSeek-V3, DeepSeek-R1, GPT series, Llama4), we run evaluations via APIs from TogetherAI [^4] https://www.together.ai and OpenAI [^5] https://platform.openai.com, using versions GPT-o1-2024-12-17, GPT-o3-mini-2025-01-31, and GPT-4o-2024-08-06. All open-sourced models up to 72B parameters were conducted with vLLM [15] framework on 8 A100 GPUs (80GB each), with a max generation length of 1,024 tokens and temperature setting of 0. Given the extreme length of some evaluated dataset (e.g., Dm-Complong with a average length of 40K tokens), we use a zero-shot setting, aim to evaluating LLMs’ capabilities for financial reasoning without any few-shot demonstrations.
Appendix E Details of Evaluated Models
To comprehensively evaluate the performance of LLMs in financial tasks, particularly those with reasoning enhancements, we select a diverse set of 29 LLMs. This selection includes both general-purpose, and finance-specific models, ranging from smaller models (8B) to large-scale models (70B), and extending to ultra-large models such as DeepSeek-R1 (671B), GPT-4o, and GPT-o1. The evaluated models are summarized as follows: GPT-4o [11]: OpenAI’s flagship model with multimodal capabilities and improved real-time reasoning. GPT-o1 [12]: A reasoning-optimized variant in the GPT series with strong step-by-step problem-solving ability. GPT-o3-mini: A compact version of GPT-o3, optimized for efficiency with robust language understanding. DeepSeek-V3 [17]: A 671B Mixture-of-Experts (MoE) model with 37B active parameters per token, designed for efficient reasoning. DeepSeek-R1 [10]: A first-generation reasoning model with competitive performance on math, code, and reasoning tasks. Qwen series [40]: A strong multilingual model with advanced instruction tuning. LLaMA-series [8]: Widely adopted open-source models (8B–70B) as well as the latest MOE models with strong instruction-tuned variants used in our experiments. DeepSeek-R1-Distill [10]: Distilled models from DeepSeek-R1, fine-tuned using LLaMA and Qwen backbones to balance performance and efficiency. Financial-Specific Reasoning Models: This category includes models such as Fin-R1-7B [19] and Dianjin-R1-7B [48], both of which leverage financial reasoning datasets and employ fine-tuning and reinforcement learning techniques to optimize LLM performance in financial reasoning tasks. Other reasoning models: They include S1-32B [22], Limo [43]. These two models are designed for general reasoning tasks, with limited but high quality COT data.
More details of the evaluated models are presented in Table 5. It also highlights the diverse training strategies employed across the evaluated models. Reinforcement learning has been extensively utilized to enhance reasoning capabilities, with common methods including PPO, DPO, and GRPO. Among these, GRPO is the most widely adopted, as seen in models such as DeepSeek-R1, Qwen2.5-72B-Instruct, and FinR1-7B. This widespread use of GRPO underscores its effectiveness in refining reasoning abilities, particularly in complex, multi-step tasks involving financial documents.
Table 5: Summary of evaluated LLMs, including their parameter sizes, reasoning capabilities, input limits, source availability, and reasoning enhancement strategies. MCTS refers to Monte Carlo Tree Search. CoT refers to chain-of-thought reasoning. TIR means tool-integrated reasoning.
Model Name Parameters Training Window Size Close/Open Source Reasoning Training Data
GPT-4o - - 128k Closed -
GPT-o1 - RL with long CoT 128k Closed Public and proprietary data (Human-annotated CoT, MCTS-assisted Synthetic data)
GPT-o3-mini - - 128k Closed Public and proprietary data
DeepSeek-V3 671B GRPO 128k Open -
DeepSeek-R1 671B GRPO 128k Open Cold-start data generation, post-processing data
GPT-4.5 - - 128k Closed -
Llama4-Scout 17B*16E SFT & DPO 128k Open -
Qwen2.5-72B-Instruct 72B DPO & GRPO 128k Open -
Qwen2.5-72B-Instruct-Math 72B DPO & GRPO 128k Open Synthetic data from Qwen, high-quality mathematical data, CoT, TIR
DeepSeek-R1-Distill-Llama-70B 70B GRPO 128k Open Distilled from R1
Llama3-70B-Instruct 70B SFT & DPO 8k Open -
Llama3.1-70B-Instruct 70B SFT & DPO 128k Open -
Llama3.3-70B-Instruct 70B SFT & DPO 128k Open -
Qwen2.5-32B-Instruct 32B DPO & GRPO 128k Open -
Qwen3-32B 32B RL with long CoT 128k Open -
DeepSeek-R1-Distill-Qwen-32B 32B GRPO 128k Open Distilled from R1
Limo 32B SFT 128k Open Public data with Qwen filteration, with CoT generated from Deepseek
S1-32B 32B SFT 128k Open Public data and original data, filtered by quality, with CoT generated from Gemini
Qwen/QwQ-32B 32B PPO 128k Open -
Qwen2.5-14B-Instruct 14B DPO & GRPO 128k Open -
Qwen3-14B 14B RL with long CoT 128k Open -
DeepSeek-R1-Distill-Qwen-14B 14B GRPO 128k Open Distilled from R1
DeepSeek-R1-Distill-Llama-8B 8B GRPO 128k Open Distilled from R1
Llama3-8B-Instruct 8B SFT & DPO 8k Open -
Llama3.1-8B-Instruct 8B SFT & DPO 128k Open -
Qwen2.5-7B-Instruct 7B DPO & GRPO 128k Open -
Qwen3-8B 8B - 128k Open -
FinR1-7B 7B SFT & GRPO 128k Open 60k CoT data for financial tasks
Dianjin-R1-7B 7B SFT & GRPO 128k Open 31k CoT data for financial tasks
@@ Appendix E Details of Evaluated Models
Appendix F Error Analysis
@@
Appendix G Prompt
@@
Limitations
Future Direction
Ethics Statement
Potential impacts
The proposed work presents both positive and negative societal implications. The development of advanced financial reasoning models has the potential to enhance financial analysis by providing more accurate predictions and robust risk assessments, thereby supporting informed decision-making in areas such as investment strategies and regulatory compliance. However, these models may also pose risks if misused, such as reinforcing existing biases in financial datasets or facilitating manipulative financial practices. To mitigate such risks, we incorporate safeguards that enforce controlled access to the models, implement guidelines for responsible use, and maintain transparency in model outputs. Additionally, we emphasize the importance of adhering to ethical guidelines throughout the development and deployment process to minimize unintended societal consequences.